# OTSeg: Multi-prompt Sinkhorn Attention for Zero-Shot Semantic Segmentation

### v1

@@ Abstract
The recent success of CLIP has demonstrated promising results in zero-shot semantic segmentation by transferring muiltimodal knowledge to pixel-level classification. However, leveraging pre-trained CLIP knowledge to closely align text embeddings with pixel embeddings still has limitations in existing approaches. To address this issue, we propose OTSeg, a novel multimodal attention mechanism aimed at enhancing the potential of multiple text prompts for matching associated pixel embeddings. We first propose Multi-Prompts Sinkhorn (MPS) based on the Optimal Transport (OT) algorithm, which leads multiple text prompts to selectively focus on various semantic features within image pixels. Moreover, inspired by the success of Sinkformers in unimodal settings, we introduce the extension of MPS, called Multi-Prompts Sinkhorn Attention (MPSA), which effectively replaces cross-attention mechanisms within Transformer framework in multimodal settings. Through extensive experiments, we demonstrate that OTSeg achieves state-of-the-art (SOTA) performance with significant gains on Zero-Shot Semantic Segmentation (ZS3) tasks across three benchmark datasets. We will release our source code at https://github.com/OTSeg.git.
@@
**footnotetext: These authors contributed equally to this work
@@ 2.2 Optimal Transport
Optimal transport (OT) is a general mathematical framework to evaluate correspondences between two distributions. Thanks to the luminous property of distribution matching, the optimal transport has received great attention and proven its generalization capability in various computer vision tasks, such as domain adaptation [12], semantic correspondence problem [19], graph matching [31, 32], and cross-domain alignment [6], etc. Among various methods, Sinkhorn algorithm can efficiently solve the OT problem through entropy-regularization [8], and it can be directly applied to deep learning frameworks thanks to the extension of Envelop Theorem [24]. In the context of computer vision tasks, prompt learning with optimal transport (PLOT) [5] optimizes the OT distance to align visual and text features by the Sinkhorn given trainable multiple text prompts for few-shot image-level prediction tasks. The most related work to ours is Sinkformer[28], where SoftMax layers within self-attention transformer is replaced by Sinkhorn algorithm, resulting in enhanced accuracy in each vision and natural language processing task. Whereas, our proposed OTSeg applies Sinkhorn algorithm for boosting cross-attention mechanism for multimodal tasks.
@@ 3.1 Optimal Transport Problem and Sinkhorn
Optimal transport aims to minimize the transport distance between two probability distributions. In this paper, we only consider discrete distribution which is closely related to our framework. We assume discrete empirical distributions $\mathbf{\mu}$ and $\mathbf{\nu}$ that are defined on probability space $\mathcal{F},\mathcal{G}\in\Omega$, respectively, as follows:
$\mathbf{\mu}=\sum^{M}_{i=1}p_{i}\delta_{f_{i}},\quad\mathbf{\nu}=\sum^{N}_{j=1}q_{j}\delta_{g_{i}},$ (1)
where $\delta_{f}$ and $\delta_{g}$ denote Dirac functions centered on ${f}$ and ${g}$, respectively, $M$ and $N$ denote the dimension of the empirical distribution. The weight vectors $\boldsymbol{p}=\{p_{i}\}^{M}_{i=1}$ and $\boldsymbol{q}=\{q_{i}\}^{N}_{j=1}$ belong to the $M$ and $N$-dimensional simplex, respectively, i.e., $\sum^{M}_{i=1}p_{i}=1$, $\sum^{N}_{j=1}q_{j}=1$. The discrete optimal transport problem can be then formulated as:
@@
However, solving the problem (Eq. 2) costs $O(n^{3}\log n)$-complexity ($n$ proportional to $M$ and $N$), which is time-consuming. This issue can be efficiently solved by the entropy-regularization of the objective through the Sinkhorn-Knopp (or simply Sinkhorn) algorithm [8]. In Sinkhorn algorithm, the optimization problem is reformulated as:
@@
where $H(\mathbf{T})$ = $\sum_{ij}\mathbf{T}_{ij}\log\mathbf{T}_{ij}$ and $\epsilon>0$ is the regularization parameter. For the problem (Eq. 3), we have an optimization solution with fewer iterations as follows:
$\texttt{Sinkhorn}\left(\frac{\mathbb{C}}{\epsilon}\right)=\mathbf{T}^{\ast}=\text{diag}(\mathbf{a}^{t})\exp(-\mathbb{C}/\epsilon)\text{diag}(\mathbf{b}^{t})$ (4)
where $t$ is the iteration and $\mathbf{a}^{t}=\mathbf{\mu}/\exp(-\mathbb{C}/\epsilon)\mathbf{b}^{t-1}$ and $\mathbf{b}^{t}=\mathbf{\nu}/\exp(-\mathbb{C}/\epsilon)\mathbf{a}^{t}$, with the initialization on $\mathbf{b}^{0}=\mathbf{1}$. To stabilize the iterative computations, we adopt the log scaling version of Sinkhorn optimization [29].
@@ 3.2 Self-Attention and Sinkformer
Sinkformer $\text{Sinkformer Attention}({\mathbf{Q}},{\mathbf{K}},{\mathbf{V}})=\texttt{Sinkhorn}\left(\frac{{\mathbf{Q}}{\mathbf{K}}^{\top}}{\sqrt{d}}\right){\mathbf{V}},$ (7)
@@ 4.2 Optimal Transformer for Zero-Shot Segmentation
where the superscript ${}^{\top}$ refers to the transpose operation, both ${\mathbf{\tilde{h}}_{\text{txt}}}$ and ${\mathbf{h}_{\text{img}}}$ are $\mathcal{L}_{2}$ normalized along the embedding dimension, and the score map ${\mathbf{S}}\in\mathbb{R}^{HW\times KN}$ undergoes further refinement.
@@
where $\;\mathbf{T}^{*}=\texttt{Sinkhorn}\left(\frac{{\mathbf{C}}}{\epsilon}\right),\;$ (10)
@@ 5.2 Implementation Details
We implemented the proposed method using the open-source toolbox MMSegmentation [7]***https://github.com/open-mmlab/mmsegmentation. The algorithm was executed on up to 8 NVIDIA A100 GPUs with a batch size of 16. We utilized the pre-trained CLIP ViT-B/16 model***https://github.com/openai/CLIP as the backbone VLM for all experiments. We fine-tuned the CLIP image encoder module by employing visual prompt tuning (VPT) approaches [15], while keeping the CLIP text encoder module frozen. The number of multiple text prompts was set to $N=6$ for VOC 2012 and $N=8$ for PASCAL Context and COCO-Stuff164K datasets. For the image decoder, we adopted a lightweight transformer consisting of three layers, with the original multi-head attention replaced by our MPSA. For the inductive settings, the total training iterations were 20K for VOC 2012, 40K for PASCAL Context, and 80K for COCO-Stuff164K. In the transductive settings, the model was trained on seen classes in the first half of training iterations and then applied self-training by generating pseudo-labels in the remaining iterations, following previous approaches [36, 37]. The optimizer was set to AdamW with a specific training schedule for each dataset.
@@ 5.3 Comparison in Zero-Shot Settings
Quantitative zero-shot segmentation results are presented in Tab. 1. We find that our proposed OTSeg outperforms the previous SOTA in both inductive and transductive settings. It implies that our proposed MPSA module decoder effectively enhances segmentation performance. Furthermore, we observe that our OTSeg+ achieves the best performance on all datasets and demonstrates the effectiveness of our ensemble strategy. Fig. 4 shows the qualitative zero-shot segmentation performance of our OTSeg+ and other previous approaches. OTSeg+ provides the most promising performance for both seen and unseen class objects compared to the previous SOTA methods. More visual results are provided in Appendix 0.I.
@@
OTSeg 13.8 ${61.9}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}-0.8}}$ ${23.6}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}-2.0}}$
OTSeg+ 13.8 ${61.9}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}-0.8}}$ ${22.5}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}-3.1}}$
@@ 7 Conclusion
Potential negative impact In our paper, we have taken measures to address potential privacy concerns related to visual results involving individuals. To safeguard privacy, we have anonymized all individuals in our figures by covering their faces.
@@ Appendix 0.A Algorithm of MPS
As discussed in Sec. 3.1, we adopt the log-scaling version of Sinkhorn approaches for our proposed Multiple Prompt Sinkhorn algorithm as follow:
@@ Appendix 0.D Definition of Harmonic mean IoU
hIoU=2*mIoU (S) * mIoU (U)mIoU (S) + mIoU ($\text{hIoU}=\frac{2*\text{mIoU (S) * mIoU (U)}}{\text{mIoU (S) + mIoU (U)}}$ (22)
@@ Appendix 0.F Implementation Detail
We further declare the implementation detail for our work. Input image resolution is set as 480$\times$480 for PASCAL Context, and 512$\times$512 for the rest of the datasets. We choose the lightest training schedule, which is 20K / 40K / 80K for VOC 2012 / PASCAL Context / COCO-Stuff-164K. The balance factor $\lambda$ for the score map in Eq. 17 is set to 0.5.
Appendix 0.G Further Component Analysis
In this section, we conduct additional component analysis by varying the conditions comprising our proposed method. Balance Factor In OTSeg+, we generate the final prediction by ensembling the decoder path and score map path as indicated in Eq. 17. Tab. S1 (a) represents the analysis of the balance factor $\lambda$. In inductive settings, the effect of the balance factor is significant compared to transductive settings. We empirically determine our default configuration of the balance factor as 0.5.
@@ Appendix 0.G Further Component Analysis
Appendix 0.H Further Score Map Comparison with or without MPSA

### v2

@@ Abstract
The recent success of CLIP has demonstrated promising results in zero-shot semantic segmentation by transferring muiltimodal knowledge to pixel-level classification. However, leveraging pre-trained CLIP knowledge to closely align text embeddings with pixel embeddings still has limitations in existing approaches. To address this issue, we propose OTSeg, a novel multimodal attention mechanism aimed at enhancing the potential of multiple text prompts for matching associated pixel embeddings. We first propose Multi-Prompts Sinkhorn (MPS) based on the Optimal Transport (OT) algorithm, which leads multiple text prompts to selectively focus on various semantic features within image pixels. Moreover, inspired by the success of Sinkformers in unimodal settings, we introduce the extension of MPS, called Multi-Prompts Sinkhorn Attention (MPSA), which effectively replaces cross-attention mechanisms within Transformer framework in multimodal settings. Through extensive experiments, we demonstrate that OTSeg achieves state-of-the-art (SOTA) performance with significant gains on Zero-Shot Semantic Segmentation (ZS3) tasks across three benchmark datasets. We release our source code at https://github.com/cubeyoung/OTSeg.
@@ 2.2 Optimal Transport
Optimal transport (OT) is a general mathematical framework to evaluate correspondences between two distributions. Thanks to the luminous property of distribution matching, the optimal transport has received great attention and proven its generalization capability in various computer vision tasks, such as domain adaptation [12], semantic correspondence problem [19], graph matching [31, 32], and cross-domain alignment [6], etc. Among various methods, Sinkhorn algorithm can efficiently solve the OT problem through entropy-regularization [8], and it can be directly applied to deep learning frameworks thanks to the extension of Envelop Theorem [24]. In the context of computer vision tasks, prompt learning with optimal transport (PLOT) [5] optimizes the OT distance to align visual and text features by the Sinkhorn given trainable multiple text prompts for few-shot image-level prediction tasks. The most related work to ours is Sinkformer[28], where SoftMax layers within self-attention transformer is replaced by Sinkhorn algorithm, resulting in enhanced accuracy in each vision and natural language processing task. The Sinkformer has inspired us to propose the Sinkhorn algorithm as an ideal fit for multimodal alignment, leading us to apply OT to further boost the performance of the cross-attention mechanism for ZS3 tasks.
@@ 3.1 Optimal Transport Problem and Sinkhorn
Optimal transport aims to minimize the transport distance between two probability distributions. In this paper, we only consider discrete distribution which is closely related to our framework. We assume the feature vector $F,G$ are defined as $F=\{\boldsymbol{f}_{i}\}_{i=1}^{M}$ and $G=\{\boldsymbol{g}_{j}\}_{j=1}^{N}$ and discrete empirical distributions $\mathbf{u}$ and $\mathbf{v}$ that are defined on probability space $\mathcal{F},\mathcal{G}\in\Omega$, respectively, as follows:
$\mathbf{u}=\sum^{M}_{i=1}\mu_{i}\delta_{\boldsymbol{f}_{i}},\quad\mathbf{v}=\sum^{N}_{j=1}\nu_{j}\delta_{\boldsymbol{g}_{j}},$ (1)
where $\delta_{\boldsymbol{f}}$ and $\delta_{\boldsymbol{g}}$ denote Dirac functions centered on $\boldsymbol{f}$ and $\boldsymbol{g}$, respectively, $M$ and $N$ denote the dimension of the empirical distribution. The weight vectors $\boldsymbol{\mu}=\{\mu_{i}\}^{M}_{i=1}$ and $\boldsymbol{\nu}=\{\nu_{j}\}^{N}_{j=1}$ belong to the $M$ and $N$-dimensional simplex, respectively, i.e., $\sum^{M}_{i=1}\mu_{i}=1$, $\sum^{N}_{j=1}\nu_{j}=1$. The discrete optimal transport problem can be then formulated as:
@@
However, solving the problem Eq. 2 costs $O(n^{3}\log n)$-complexity ($n$ proportional to $M$ and $N$), which is time-consuming. This issue can be efficiently solved by the entropy-regularization of the objective through the Sinkhorn-Knopp (or simply Sinkhorn) algorithm [8]. In Sinkhorn algorithm, the optimization problem is reformulated as:
@@
where $H(\mathbf{T})$ = $\sum_{ij}\mathbf{T}_{ij}\log\mathbf{T}_{ij}$ and $\epsilon>0$ is the regularization parameter. For the problem Eq. 3, we have an optimization solution when $t\rightarrow\infty$ as follow:
$\mathbf{T}^{\ast}=\text{diag}(\mathbf{a}^{t})\exp(-\mathbb{C}/\epsilon)\text{diag}(\mathbf{b}^{t})$ (4)
where $t$ is the iteration and $\mathbf{a}^{t}=\boldsymbol{\mu}/\exp(-\mathbb{C}/\epsilon)\mathbf{b}^{t-1}$ and $\mathbf{b}^{t}=\boldsymbol{\nu}/\exp(-\mathbb{C}/\epsilon)\mathbf{a}^{t}$, with the initialization on $\mathbf{b}^{0}=\mathbf{1}$. To stabilize the iterative computations, we adopt the log scaling version of Sinkhorn optimization [29].
@@ 3.2 Self-Attention and Sinkformer
$\text{Sinkformer Attention}({\mathbf{Q}},{\mathbf{K}},{\mathbf{V}})=\texttt{Sinkhorn}\left(\frac{{\mathbf{C}}}{\sqrt{d}}\right){\mathbf{V}},\;{\mathbf{C}}=\boldsymbol{1}-{\mathbf{Q}}{\mathbf{K}}^{\top}$ (7)
@@ 4.2 Optimal Transformer for Zero-Shot Segmentation
where the superscript refers to the transpose operation, both ${\mathbf{\tilde{h}}_{\text{txt}}}$ and ${\mathbf{h}_{\text{img}}}$ are $\mathcal{L}_{2}$ normalized along the embedding dimension, and the score map ${\mathbf{S}}\in\mathbb{R}^{HW\times KN}$ undergoes further refinement.
@@
where $\;\mathbf{T}^{*}=\texttt{Sinkhorn}\left(\frac{{\mathbf{C}}}{\epsilon}\right),\;\mathbf{C}:=\mathbf{1-S}\;$ (10)
@@ 5.2 Implementation Details
We implemented the proposed method using the open-source toolbox MMSegmentation [7]. The algorithm was executed on up to 8 NVIDIA A100 GPUs with a batch size of 16. We utilized the pre-trained CLIP ViT-B/16 model [^1] https://github.com/openai/CLIP as the backbone VLM for all experiments. We fine-tuned the CLIP image encoder module by employing visual prompt tuning (VPT) approaches [15], while keeping the CLIP text encoder module frozen. The number of multiple text prompts was set to $N=6$ for VOC 2012 and $N=8$ for PASCAL Context and COCO-Stuff164K datasets. For the image decoder, we adopted a lightweight transformer consisting of three layers, with the original multi-head attention replaced by our MPSA. The optimizer was set to AdamW with a specific training schedule for each dataset.
@@ 5.3 Comparison in Zero-Shot Settings
Quantitative zero-shot segmentation results are presented in Tab. 1. We find that our proposed OTSeg outperforms the previous SOTA in both inductive and transductive settings. It implies that our proposed MPSA module decoder effectively enhances segmentation performance. Furthermore, we observe that our OTSeg+ achieves the best performance on all datasets and demonstrates the effectiveness of our ensemble strategy. Fig. 4 shows the qualitative zero-shot segmentation performance of our OTSeg+ and other previous approaches. OTSeg+ provides the most promising performance for both seen and unseen class objects compared to the previous SOTA methods. More visual results are provided in Appendix 0.K.
@@
OTSeg 13.8 ${61.9}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.8}}$ ${23.6}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-2.0}}$
OTSeg+ 13.8 ${61.9}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.8}}$ ${22.5}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-3.1}}$
@@ 7 Conclusion
Acknowledgments This work was supported by Institute for Information &\&& communications Technology Promotion(IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II190075) Artificial Intelligence Graduate School Program(KAIST), and also supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2024-00345854), (RS-2024-00336454), (RS-2023-00262527), and also supported by Field-oriented Technology Development Project for Customs Administration funded by the Korea government through the National Research Foundation (NRF) of Korea under Grant NRF2021M3I1A1097910, and also Administration by Culture, Sports and Tourism R&\&&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2023.
@@ Appendix 0.A Algorithm of MPS
As discussed in Sec. 3.1, we adopt the log-scaling version of Sinkhorn approaches for our proposed Multiple Prompt Sinkhorn (MPS) algorithm as follow:
@@ Appendix 0.D Definition of Harmonic mean IoU
hIoU=2mIoU (S) * mIoU (U)mIoU (S) + mIoU ($\text{hIoU}=\frac{2*\text{mIoU (S) * mIoU (U)}}{\text{mIoU (S) + mIoU (U)}}$ (22)
@@ Appendix 0.F Implementation Detail
We further declare the implementation detail for our work. Input image resolution is set as 480$\times$480 for PASCAL Context, and 512$\times$512 for the rest of the datasets. For training, we choose the lightest training schedule. For the inductive settings, the total training iterations were 20K for VOC 2012, 40K for PASCAL Context, and 80K for COCO-Stuff164K. For the transductive settings, the model was trained on seen classes in the first half of training iterations and then applied self-training by generating pseudo-labels in the remaining iterations, following previous approaches [37, 38]. The balance factor $\lambda$ for the score map in Eq. 17 is set to 0.5.
Appendix 0.G Further Comparison Studies
To further validate the generalizability of our proposed method in the open-vocabulary segmentation (OVS) setting, we compared our OTSeg and OTSeg+ with FreeSeg [25] and SAN [34]. It’s noteworthy that OTSeg is designed for ZS3, utilizing a partial dataset of 156 classes, whereas FreeSeg and SAN use the entire COCO-Stuff 171 classes dataset tailored for OVS. Despite differing goals and settings, our proposed OTSeg variants demonstrate comparable results in cross-dataset settings, as shown in Table S1, validating their effectiveness.
Table S1: Cross-data comparison with open-vocabulary segmentation methods.
Category Method Source ADE20K PC59 VOC
Zero-shot segmentation (ZS3) OTSeg COCO-156 21.9 52.9 94.2
OTSeg+ 21.1 53.4 94.4
Open-vocabulary segmentation (OVS) Freeseg [25] COCO-171 17.9 34.4 85.6
SAN[34] 27.5 53.8 94.1
Appendix 0.H Robustness of OTSeg with Multiple Seeds
To evaluate the robustness of our proposed Multiple Prompt Sinkhorn (MPS) algorithm, we repeatedly trained both OTSeg and OTSeg+ with multiple seeds and compared the results on the VOC dataset. Table S2 confirms the robustness and consistent performance of our proposed method across various seeds.
Table S2: Quantitative results with multiple seeds on VOC dataset and comparison in cross-data settings with further approaches.
Settings Method Seed Mean $\pm$ STD
#1 #2 #3 #4
Inductive OTSeg 84.5 84.9 84.3 84.0 84.4 $\pm$ 0.4
OTSeg+ 87.1 88.0 88.0 86.5 87.4 $\pm$ 0.7
Transductive OTSeg 94.2 94.6 94.2 94.3 94.4 $\pm$ 0.2
OTSeg+ 94.3 94.5 94.2 94.3 94.3 $\pm$ 0.1
Appendix 0.I Further Component Analysis
In this section, we conduct additional component analysis by varying the conditions comprising our proposed method.
Balance Factor In OTSeg+, we generate the final prediction by ensembling the decoder path and score map path as indicated in Eq. 17. Tab. S3 (a) represents the analysis of the balance factor $\lambda$. In inductive settings, the effect of the balance factor is significant compared to transductive settings. We empirically determine our default configuration of the balance factor as 0.5.
@@ Appendix 0.I Further Component Analysis
Appendix 0.J Analysis on Sinkhorn Convergence Comparison Studies
Figure S1: Sinkhorn convergence.
We further analyzed the convergence of the Sinkhorn algorithm and its effect on performance. We established a defined threshold and applied the Sinkhorn algorithm iteratively until the total errors dropped below this threshold. In Figure S1, we report that the Sinkhorn algorithm converged in 6 iterations for the score map and 14 iterations for the OT-adapted decoder part on the VOC test dataset. Additionally, we observed that the convergence of the Sinkhorn algorithm around 15K training iterations led to the best performance.
Appendix 0.K Additional Visual Results
Appendix 0.L Further Score Map Comparison with or without MPSA